The sum of a number and 8 times its reciprocal is 6. Find the number(s).
Answer by RAY100(1637)

Let n= #,,,,,,,reciprocal = 1/n
.
n+ 8(1/n) = 6,,,given,,,,,,multiply thru by n
.
n^2 +8 = 6n
.
n^2 -6n +8 = 0,,,,,,factor
.
(n-4)(n-2) = 0
.
n= 2, 4
.
.
check,,,,(2),,,,,, 2 + 8*1/2 =6,,,,,ok
.
(4),,,,,,4 +8*1/4 = 6,,,,,,ok
.